# Theory of Valley Splitting in Si/SiGe Spin-Qubits: Interplay of Strain, Resonances and Random Alloy Disorder

### v2

@@ IV Discussion
The theoretical model developed in Sec. II of this work combines a number of previously existing modeling concepts into a unified framework. In this way, a comprehensive description of a number of important effects contributing to the valley splitting in engineered heterostructures, i.e., alloy-disorder, strain and resonances, has been achieved at the level of envelope function theory.
@@
The numerical results in Sec. III are consistent with previously reported findings, but also extend them in view of a systematic investigation of strain effects. For instance, significant shear strain-induced enhancements of the valley splitting even at wide interfaces were predicted. In addition, the model includes all non-trivial resonances that mediate intervalley coupling across neighboring Brillouin zones. Most notably, this includes the important resonance at $2k_{1}$, which triggers the mechanism behind the long-period wiggle well. This is the main difference with the prevailing $2k_{0}$-theory, which does not explain the long-period wiggle-well. The key accomplishments of the $2k_{0}$-theory as described in Refs. [19, 20], lies in the description of disorder-induced effects and the distinction between deterministic and disorder-dominated enhancements of valley splitting. This part of the theory has been adopted practically identically in the present model.
@@
The model presented in this paper differs in a number of aspects from the model by Feng & Joynt [31], which has been the first to provide an explanation for the mechanism behind the long-period wiggle-well in the framework of envelope function theory. The mechanism proposed there to induce a violation of the nonsymmorphic screw symmetry – which otherwise suppresses the resonance at $2k_{1}$ – is based on a randomization of the plane wave expansion coefficients of the Bloch factors. The latter are determined for statistical ensembles of disordered SiGe alloys using a combination of an empirical pseudopotential model and an extended virtual crystal approximation, whereby the alloy disorder induces the desired symmetry breaking. This approach is in principle very plausible, but a subsequent analysis shows that the corresponding statistical distribution of the valley splitting implied by this method does not lead to a Rice distribution (as observed in experiments and atomistic simulations). Instead, a significantly more skewed distribution is found, which has a significantly larger probability density at low valley splittings compared to the Rice distribution. Furthermore, in the model by Feng & Joynt, the distribution of the intervalley coupling parameter in the complex plane takes a strongly elliptical shape, which is very different from the nearly circular distribution obtained here (see Appendix C.1). Another consequence of the symmetry-breaking mechanism proposed there is that the variance of the valley splitting in the long-period wiggle-well becomes very large. Thus, the valley splitting enhancement due to the long-period wiggle-well is found to be primarily disorder-dominated and is by no means deterministic. This is reasonable when one assumes that the symmetry breaking is disorder-induced. These predictions on the characteristics of the long-period wiggle-well are in sharp contrast to those obtained in the present work, where the nonsymmorphic screw symmetry is broken via shear strain. As a consequence, the present model predicts a strong deterministic enhancement of the valley splitting in the long-period wiggle-well. The actual microscopic symmetry-breaking mechanism, however, is very likely a combination of both effects. The impact of disorder on the long-period wiggle-well might therefore require further investigation. Finally, the model by Feng & Joynt shows another pronounced resonance for a wiggle well with intermediate wave number $q=k_{0}$. In our model, we found that this is caused by a harmonic similar to the configuration described in Sec. III.2.4. For typical SiGe parameters (recall that Feng & Joynt used a conduction band offset more typical for Si-MOS), however, the small deterministic resonance is entirely obscured by the disorder-induced component. Therefore, no peak can be observed in the mean valley splitting in Fig. 4.
@@
Finally, we discuss the relation between the present work and the recent paper by Woods et al. [33]. In Ref. [33], the shear strain induced symmetry-breaking mechanism to unlock the resonance behind the long-period wiggle-well has been studied for the first time. The envelope function model presented there, which is derived from a one-dimensional tight-binding model, features an intervalley coupling matrix element with unconventional functional form that is notably different from other expressions in the literature. In the present model described in Sec. II of this work, however, the standard form of the intervalley matrix element is preserved even in the presence of strain effects. The numerical results obtained here are qualitatively in an excellent agreement with those reported by Woods et al.. In particular the joint linear dependence of the long-period wiggle-well on shear strain and Ge concentration is recovered using the empirical pseudopotential model. Furthermore, we also find good quantitative agreements in the magnitude of the obtained valley splittings. The quantitative comparison, however, requires caution, as the present model also contains disorder-induced contributions, which are omitted in Ref. [33].
@@ V Conclusions and Outlook
We have developed a comprehensive theoretical model that describes a broad range of physical mechanisms determining the valley splitting in Si/SiGe heterostructures, namely alloy-disorder, strain and resonances between valley states across neighboring Brillouin zones. The numerical results are consistent with previously known results but also offer new insights extending the state of the art. With regard to applications, the key accomplishment of the model presented here is a comprehensive characterization of the long-period wiggle-well. This includes complex phenomena such as the shear strain-induced symmetry breaking, which unlocks the corresponding resonance mechanism, as well as a quantification of the statistical fluctuations of the valley splitting caused by alloy disorder. Our model predicts a significant deterministic enhancement of the valley splitting for the long-period wiggle-well. Due to the critical role of shear strain, we conclude that strain engineering [43, 58, 33] must be further advanced to enhance the design of Si/SiGe spin-qubits.
In contrast to detailed atomistic models, the present model offers a comprehensive description of multiple physical phenomena at very low computational cost. Therefore, it is well suited to rapidly characterize the valley splittings of a vast number of different epitaxial designs. In a subsequent work, the present model will be employed for free-shape optimization of the epitaxial profile to further enhance the valley splitting beyond the simple heuristics considered here.
@@ Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689, project AA2–17). M. K. is grateful for valuable discussions with Mark Friesen, Merritt P. Losert and Lars R. Schreiber.
Appendix A Derivation of the Multi-Valley Coupled Envelope Equation Model
In this section, we provide a derivation of the coupled envelope equations (1) loosely following Ref. [66].
@@ Appendix A Derivation of the Multi-Valley Coupled Envelope Equation Model
$H\left(\mathbf{r}\right)\Psi_{n}\left(\mathbf{r}\right)=E_{n}\Psi_{n}\left(\mathbf{r}\right).$ (41)
The full Hamiltonian reads
$H\left(\mathbf{r}\right)=H_{\mathrm{cr}}\left(\mathbf{r}\right)+U\left(\mathbf{r}\right),$ (42)
which includes the lattice-periodic crystal Hamiltonian
$H_{\mathrm{cr}}\left(\mathbf{r}\right)=-\frac{\hbar^{2}}{2m_{0}}\nabla^{2}+V_{0}\left(\mathbf{r}\right)$ (43)
@@
and the confinement potential $U\left(\mathbf{r}\right)$ described in Eq. (4). Here, $m_{0}$ is the vacuum electron mass and $V_{0}\left(\mathbf{r}\right)$ is the lattice-periodic potential of the crystal. The crystal potential is invariant under translations with lattice vectors $V_{0}\left(\mathbf{r}\right)\equiv V_{0}\left(\mathbf{r}+\mathbf{R}_{j}\right)$. We assume that $V_{0}$ describes solely the perfectly periodic Si crystal, whereas the effects of Ge in both the QW and the barrier are described by the heterostructure potential that is part of $U\left(\mathbf{r}\right)$, see Eq. (4). In this way, double-counting of Ge atoms is avoided and alloy disorder effects are separated from the idealized band structure computation.
The translation-invariant lattice-periodic part is solved by Bloch functions that are indexed by a band index $n$ and the wave vector $\mathbf{k}$ as
$H_{\mathrm{cr}}\left(\mathbf{r}\right)\phi_{n,\mathbf{k}}\left(\mathbf{r}\right)$ $=E_{n,\mathbf{k}}\phi_{n,\mathbf{k}}\left(\mathbf{r}\right).$ (44)
The Bloch function reads
@@
$\phi_{n,\mathbf{k}}\left(\mathbf{r}\right)=\frac{1}{\sqrt{V}}\mathrm{e}^{i\mathbf{k}\cdot\mathbf{r}}u_{n,\mathbf{k}}\left(\mathbf{r}\right),$ (45)
where $u_{n,\mathbf{k}}\left(\mathbf{r}\right)\equiv u_{n,\mathbf{k}}\left(\mathbf{r}+\mathbf{R}_{j}\right)$ is the lattice-periodic Bloch factor and $V$ is the crystal volume. The Bloch functions form a complete orthonormal basis
@@
$\int_{V}\mathrm{d}^{3}r\,\phi_{n,\mathbf{k}}^{*}\left(\mathbf{r}\right)\phi_{n^{\prime},\mathbf{k}^{\prime}}\left(\mathbf{r}\right)=\delta_{n,n^{\prime}}\,\delta_{\mathbf{k},\mathbf{k}^{\prime}}$ (46)
and the Bloch factors obey an orthogonality relation for identical wave vectors
$\frac{1}{\Omega_{p}}\int_{\Omega_{p}}\mathrm{d}^{3}r\,u_{n,\mathbf{k}}^{*}\left(\mathbf{r}\right)u_{n^{\prime},\mathbf{k}}\left(\mathbf{r}\right)=\delta_{n,n^{\prime}},$ (47)
where $\Omega_{p}$ is the volume of the primitive unit cell. We assume $V=N\Omega_{p}$ with $N$ unit cells.
In the presence of the heterostructure, the translation invariance is broken such that the wave vector is not a good quantum number anymore. Therefore, we choose the ansatz
$\Psi_{n}\left(\mathbf{r}\right)=\sqrt{V}\sum_{\mathbf{k}}F_{n,\mathbf{k}}\phi_{n,\mathbf{k}}\left(\mathbf{r}\right),$ (48)
where $F_{n,\mathbf{k}}$ are expansion coefficients, whose Fourier transform will lead to the envelope wave functions below. We assume normalization of the expansion coefficients
$V\sum_{\mathbf{k}}\left|F_{n,\mathbf{k}}\right|^{2}=1$ (49)
to ensure orthonormalization of the full wave function.
In order to construct the multi-valley envelope equation, the expansion coefficients have to be separated into contributions from different valleys. We assume that the coefficients are only non-zero in the vicinity of the valley minima $\mathbf{k}\approx\mathbf{k}_{\nu}$ such that, following Ref. [66], the expansion coefficient is written as
$F_{n,\mathbf{k}}=\sum_{\nu}\alpha_{\nu}F_{n,\nu}\left(\mathbf{k}-\mathbf{k}_{\nu}\right)$ (50)
where $\nu$ is the valley-index and $\alpha_{\nu}$ is a symmetry factor. We assume normalization as $\sum_{\nu}\left|\alpha_{\nu}\right|^{2}=1$.
Following Ref. [66], a Shindo–Nara-type multi-valley coupled envelope equation model is obtained by taking the mean of Eq. (41) with respect to the ansatz (48).
A.1 Lattice-Periodic Part
The lattice-periodic part of the coupled envelope equations is obtained from
$\sum_{n^{\prime}}$ $\langle\Psi_{n}|\left(\hat{H}_{\mathrm{cr}}-E_{n^{\prime}}\right)|\Psi_{n^{\prime}}\rangle=$
$=V\sum_{n^{\prime}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}F_{n,\mathbf{k}}^{*}F_{n^{\prime},\mathbf{k}^{\prime}}\times$
@@ A.1 Lattice-Periodic Part
$=V\sum_{\mathbf{k}}\left(E_{n,\mathbf{k}}-E_{n}\right)\left|F_{n,\mathbf{k}}\right|^{2},$
where we used Eqs. (44) and (46). Using the valley-separation (50) and the parabolic band approximation
$E_{n,\mathbf{k}_{\nu}+\mathbf{k}}\approx E_{n,\mathbf{k}_{\nu}}+\frac{\hbar^{2}}{2}\mathbf{k}\cdot m_{n,\mathbf{k}_{\nu}}^{-1}\mathbf{k},$ (51)
one arrives at

### v3

@@ IV Discussion
The theoretical model developed in this work combines a number of previously existing modeling concepts into a unified framework. In this way, a comprehensive description of a number of important effects contributing to the valley splitting in engineered heterostructures, i.e., alloy-disorder, strain and resonances, has been achieved at the level of envelope function theory.
@@
The numerical results in Sec. III are consistent with previously reported findings, but also extend them by systematically investigating the effects of strain. For instance, significant shear strain-induced enhancements of the valley splitting even at wide interfaces were predicted. In addition, the model includes all non-trivial resonances that mediate intervalley coupling across neighboring Brillouin zones. Most notably, this includes the important resonance at $2k_{1}$, which triggers the mechanism behind the long-period wiggle well. This is the main difference with the prevailing $2k_{0}$-theory, which does not explain the long-period wiggle-well. The key accomplishments of the $2k_{0}$-theory as described in Refs. [19, 20], lies in the description of disorder-induced effects and the distinction between deterministic and disorder-dominated enhancements of valley splitting. This part of the theory has been adopted practically identically in the present model.
The results reported in this paper depend to a minor degree on the chosen pseudopotential model. In addition to the model by Rieger & Vogl [54], we have also tested the model by Fischetti & Laux [53] and obtained qualitatively similar results with slight quantitative differences. A detailed comparison is beyond the scope of this paper.
@@
The present model differs in a number of aspects from the model by Feng & Joynt [31], which has been the first to provide an explanation for the mechanism behind the long-period wiggle-well in the framework of envelope function theory. The mechanism proposed there to induce a violation of the nonsymmorphic screw symmetry—which otherwise suppresses the resonance at $2k_{1}$—is based on a randomization of the plane wave expansion coefficients of the Bloch factors. The latter are determined for statistical ensembles of disordered SiGe alloys using a combination of an empirical pseudopotential model and an extended virtual crystal approximation, whereby the alloy disorder induces the desired symmetry breaking. This approach is in principle very plausible, but a subsequent analysis shows that the corresponding statistical distribution of the valley splitting implied by this method does not lead to a Rice distribution (as observed in experiments and atomistic simulations). Instead, a significantly more skewed distribution is found, which has an increased probability density at low valley splittings compared to the Rice distribution. Furthermore, in the model by Feng & Joynt, the distribution of the intervalley coupling parameter in the complex plane takes a strongly elliptical shape, which is very different from the nearly circular distribution obtained here (see Appendix D.1). Another consequence of the symmetry-breaking mechanism proposed there is that the variance of the valley splitting in the long-period wiggle-well becomes very large. Thus, the valley splitting enhancement due to the long-period wiggle-well is found to be primarily disorder-dominated and is by no means deterministic. This is reasonable when one assumes that the symmetry breaking is disorder-induced. These predictions on the characteristics of the long-period wiggle-well are in sharp contrast to those obtained in the present work, where the nonsymmorphic symmetry is broken via shear strain. As a consequence, the present model predicts a strong deterministic enhancement of the valley splitting in the long-period wiggle-well. The actual microscopic symmetry-breaking mechanism, however, is very likely a combination of both effects. The impact of disorder on the long-period wiggle-well might therefore require further investigation. Finally, the model by Feng & Joynt shows another pronounced resonance for a wiggle well with intermediate wave number $q=k_{0}$. In our model, we found that this is caused by a harmonic similar to the configuration described in Sec. III.2.4. For typical SiGe parameters (recall that Feng & Joynt used a conduction band offset more typical for Si-MOS), however, the small deterministic resonance is entirely obscured by the disorder-induced component. Thus, no peak at $q=k_{0}$ can be observed in the mean valley splitting in Fig. 4.
@@
Finally, we discuss the relation between the present work and the paper by Woods et al. [33]. In Ref. [33], the shear strain induced symmetry-breaking mechanism to unlock the resonance behind the long-period wiggle-well has been studied for the first time. The envelope function model presented there, which is derived from a one-dimensional tight-binding model, features an intervalley coupling matrix element with unconventional functional form that is notably different from other expressions in the literature. In our model, however, the standard form of the intervalley matrix element is preserved even in the presence of strain effects. The numerical results obtained here are qualitatively in excellent agreement with those reported by Woods et al.. In particular the joint linear dependence of the long-period wiggle-well on shear strain and Ge concentration is recovered using the EPM. Furthermore, we also find good quantitative agreements in the magnitude of the obtained valley splittings. The quantitative comparison, however, requires caution, as the present model also contains disorder-induced contributions, which are omitted in Ref. [33].
@@ V Conclusions and Outlook
We have developed a comprehensive theoretical model that describes a broad range of physical mechanisms determining the valley splitting in Si/SiGe heterostructures, namely alloy-disorder, strain and resonances between valley states across neighboring Brillouin zones. The numerical results are consistent with previously known results but also offer new insights extending the state of the art. The key accomplishment of our model is the explanation of the long-period wiggle-well in the framework of envelope function theory as a result of shear strain-induced symmetry breaking, which unlocks the corresponding resonance mechanism. Due to the critical role of shear strain, we conclude that strain engineering [44, 67, 60, 33] must be further advanced to enhance the design of Si/SiGe spin-qubits.
In contrast to detailed atomistic models, the present model offers a comprehensive description of multiple physical phenomena at very low computational cost. Therefore, it is well suited to rapidly characterize the valley splitting in different epitaxial designs. In a subsequent work, the present model will be employed for free-shape optimization of the epitaxial profile to further enhance the valley splitting beyond the heuristic strategies considered here.
Code Availability
MATLAB code to reproduce the simulation results and figures of this work is available on GitHub [68].
@@ Acknowledgements.
This work was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany’s Excellence Strategy – The Berlin Mathematics Research Center MATH+ (EXC-2046/1, project ID: 390685689, project AA2–17). M. K. is grateful for valuable discussions with Mark Friesen, Merritt P. Losert, Lars R. Schreiber and Thomas Koprucki.
Appendix A Multi-Valley Coupled Envelope Equation Model
In this section, we provide a derivation of the coupled envelope equations (1) based on a Burt–Foreman type envelope function theory for multi-valley semiconductors.
A.1 Microscopic Schrödinger Equation
@@
$\left(-\frac{\hbar^{2}}{2m_{0}}\nabla^{2}+V\left(\mathbf{r}\right)+U\left(\mathbf{r}\right)\right)\Psi\left(\mathbf{r}\right)=E\Psi\left(\mathbf{r}\right),$ (40)
@@
where $m_{0}$ is the vacuum electron mass, $V\left(\mathbf{r}\right)$ is a lattice-periodic potential and $U\left(\mathbf{r}\right)$ is a mesoscopic confinement potential. The crystal potential is invariant under translations with lattice vectors $V\left(\mathbf{r}\right)\equiv V\left(\mathbf{r}+\mathbf{R}\right)$. Throughout this paper we assume that $V_{0}$ describes solely the perfectly periodic Si crystal, whereas the effects of Ge in both the QW and the barrier are described by the heterostructure potential $U_{\mathrm{het}}\left(\mathbf{r}\right)$ that is included in $U\left(\mathbf{r}\right)$, see Eq. (4). This way, double-counting of Ge atoms is avoided and alloy disorder effects are separated from the idealized band structure computation.
A.2 Multi-Valley Envelope Wave Function
The single-particle wave function for electrons in semiconductor nanostructures shows rapid spatial oscillations on the atomistic scale of the underlying crystal as well as a slowly varying envelope on the mesoscopic scale of the confinement potential. In order to obtain an effective description of the envelope wave function, the rapid atomistic features must be eliminated. There is extensive literature on the (heuristic) derivation of envelope function theories for electrons in Si-based nanostructures [69, 70, 71, 72, 73] that employ a number of approximations, most notably the assumption of a slowly varying potential. In the context of the present work, the assumption of a slowly varying potential is questionable, as we are particularly interested in the investigation of QWs with sharp interfaces and highly oscillatory wiggle wells. Burt [74, 75] and Foreman [76, 77] have developed an exact envelope function theory that avoids these assumptions, but leads to a system of coupled integro-differential equations with a nonlocal potential. The approach has been extended to multi-valley semiconductors by Klymenko et al. [78, 79] based on a decomposition of the first Brillouin zone into non-overlapping valley-specific sectors, see Fig. 6 (a). In this section we provide a derivation of the coupled envelope equations (1) based on multi-valley Burt–Foreman type envelope function theory.
Fig. 6: (a) Decomposition of the first Brillouin zone of the fcc lattice into non-overlapping valley-specific sectors $S\left(\mathbf{k}_{0}\right)$. The sectors of the two low-energy valleys are highlighted. (b) The wave vectors within the sectors are decomposed into a valley vector $\mathbf{k}_{0}$ and a deviation $\mathbf{k}$ such that $\left(\mathbf{k}_{0}+\mathbf{k}\right)\in S\left(\mathbf{k}_{0}\right)$.
Following [78, 79], the wave function is expanded as
$\Psi\left(\mathbf{r}\right)=\sum_{n}\sum_{\mathbf{k}_{0}}u_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)\mathrm{e}^{i\mathbf{k}_{0}\cdot\mathbf{r}}f_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right),$ (41)
where $n$ labels the band index, $\mathbf{k}_{0}$ are the wave vectors of the conduction band valleys, $u_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)$ is the lattice-periodic Bloch factor and $f_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)$ is the (slowly varying) envelope wave function. The Fourier expansion of the wave function reads
$\Psi\left(\mathbf{r}\right)=\sum_{\mathbf{G}}\sum_{\mathbf{k}_{0}}\sum_{\left(\mathbf{k}+\mathbf{k}_{0}\right)\in S\left(\mathbf{k}_{0}\right)}\mathrm{e}^{i\left(\mathbf{G}+\mathbf{k}_{0}+\mathbf{k}\right)\cdot\mathbf{r}}\Psi\left(\mathbf{G}+\mathbf{k}_{0}+\mathbf{k}\right),$
where each possible wave number is uniquely described by a reciprocal lattice vector $\mathbf{G}$, a valley vector $\mathbf{k}_{0}$ and a small wave number $\mathbf{k}$ such that the sum $\mathbf{k}+\mathbf{k}_{0}$ is restricted to the sector $S\left(\mathbf{k}_{0}\right)$ of the Brillouin zone, see Fig. 6 (b). The exponential factor $\exp{\left(i\mathbf{G}\cdot\mathbf{r}\right)}$ can be expanded in terms of lattice-periodic Bloch functions
$\mathrm{e}^{i\mathbf{G}\cdot\mathbf{r}}=\sum_{n}u_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)c_{n,\mathbf{k}_{0}}^{*}\left(\mathbf{G}\right),$
such that we obtain from (41) and its Fourier transform an expression for the envelope
$f_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)=$ (42)
$=\sum_{\left(\mathbf{k}+\mathbf{k}_{0}\right)\in S\left(\mathbf{k}_{0}\right)}\mathrm{e}^{i\mathbf{k}\cdot\mathbf{r}}\sum_{\mathbf{G}}c_{n,\mathbf{k}_{0}}^{*}\left(\mathbf{G}\right)\Psi\left(\mathbf{G}+\mathbf{k}_{0}+\mathbf{k}\right).$
This function is slowly varying as its Fourier expansion is restricted to a limited set of (small) wave vectors $\mathbf{k}$ relative to the valley vector such that $\left(\mathbf{k}+\mathbf{k}_{0}\right)\in S\left(\mathbf{k}_{0}\right)$.
Following [80], the envelope equation is derived starting out from the Fourier domain representation of Schrödinger’s equation (40). Substitution of the ansatz (41) yields
$Ef_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)$ $=\left(-\frac{\hbar^{2}}{2m_{0}}\nabla^{2}-\frac{i\hbar^{2}}{m_{0}}\mathbf{k}_{0}\cdot\nabla+\frac{\hbar^{2}k_{0}^{2}}{2m_{0}}\right)f_{n,\mathbf{k}_{0}}\left(\mathbf{r}\right)$
@@ A.2 Multi-Valley Envelope Wave Function
$\phantom{=}+\frac{1}{m_{0}}\sum_{n^{\prime}}\mathbf{p}_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)\cdot\left(\hbar\mathbf{k}_{0}-i\hbar\nabla\right)f_{n,^{\prime}\mathbf{k}_{0}}\left(\mathbf{r}\right)$
$\phantom{=}+\sum_{n^{\prime}}\left(T_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)+V_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)\right)f_{n,^{\prime}\mathbf{k}_{0}}\left(\mathbf{r}\right)$
$\phantom{=}+\sum_{n^{\prime}}\sum_{\mathbf{k}_{0}^{\prime}}\int_{V}\mathrm{d}^{3}r^{\prime}\,U_{\mathbf{k}_{0},\mathbf{k}_{0}^{\prime}}^{n,n^{\prime}}\left(\mathbf{r},\mathbf{r}^{\prime}\right)f_{n^{\prime},\mathbf{k}_{0}^{\prime}}\left(\mathbf{r}^{\prime}\right)$
where the nonlocal potential energy kernel reads
$U_{\mathbf{k}_{0},\mathbf{k}_{0}^{\prime}}^{n,n^{\prime}}\left(\mathbf{r},\mathbf{r}^{\prime}\right)$ $=\mathrm{e}^{-i\left(\mathbf{k}_{0}\cdot\mathbf{r}-\mathbf{k}_{0}^{\prime}\cdot\mathbf{r}^{\prime}\right)}\times$ (43)
@@
×$\phantom{=}\times\int_{V}\mathrm{d}^{3}r^{\prime\prime}\,\Delta_{\mathbf{k}_{0}}\left(\mathbf{r}-\mathbf{r}^{\prime\prime}\right)u_{n,\mathbf{k}_{0}}^{*}\left(\mathbf{r}^{\prime\prime}\right)\times$
@@
The matrix elements of the momentum, kinetic and potential energy operator of the bulk crystal are
$\mathbf{p}_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)$ $=\frac{1}{\Omega_{p}}\int_{\Omega_{p}}\mathrm{d}^{3}r\,u_{n,\mathbf{k}_{0}}^{*}\left(\mathbf{r}\right)\left(-i\hbar\nabla u_{n^{\prime},\mathbf{k}_{0}}\left(\mathbf{r}\right)\right),$
$T_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)$ $=\frac{1}{\Omega_{p}}\int_{\Omega_{p}}\mathrm{d}^{3}r\,u_{n,\mathbf{k}_{0}}^{*}\left(\mathbf{r}\right)\left(-\frac{\hbar^{2}}{2m_{0}}\nabla^{2}u_{n^{\prime},\mathbf{k}_{0}}\left(\mathbf{r}\right)\right),$
$V_{n,n^{\prime}}\left(\mathbf{k}_{0}\right)$ $=\frac{1}{\Omega_{p}}\int_{\Omega_{p}}\mathrm{d}^{3}r\,u_{n,\mathbf{k}_{0}}^{*}\left(\mathbf{r}\right)V\left(\mathbf{r}\right)u_{n^{\prime},\mathbf{k}_{0}}\left(\mathbf{r}\right).$
The nonlocal term (43) for the mesoscopic confinement potential involves the low-pass filtered delta function